OpenStudy (anonymous):

Simplify the following expression: (5x^2 + 3x + 4) − (2x^2 + 5x − 1). If the final answer is written in the form Ax^2 + Bx + C, what is the value of A?

OpenStudy (anonymous):

(5x2 + 3x + 4) − (2x2 + 5x − 1) = 5x2 + 3x + 4 − 2x2 - 5x + 1 = 5x2 - 2x2 + 3x - 5x + 4 + 1 = 3x2 - 2x + 5. jesus

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

i dont even get wat u just did

OpenStudy (e.cociuba):

1. Answers.com is not always preferred for people when learning. 2. Don't use the Lords name in vain. 3. Describe it and help her learn the method and equation. @jonjenkins7653

OpenStudy (anonymous):

i can do what i want to jesus jesus jesus jesus

OpenStudy (anonymous):

.....

OpenStudy (anonymous):

@e.cociuba bru u can go

OpenStudy (anonymous):

how long u goin to type

OpenStudy (e.cociuba):

Okay, so we have this equation, and what we have to do with it, is simplify it: (5x^2 + 3x + 4) − (2x^2 + 5x − 1) There is no point in keeping the parenthesis, since you can combine both of them together, which gives you this: 5x^2 + 3x + 4 − 2x^2 - 5x + 1 =

OpenStudy (e.cociuba):

Does that make sense so far?

OpenStudy (anonymous):

ohh ok so what i do is combine like terms or?

OpenStudy (e.cociuba):

Yes. Next you're going to combine your like terms. Show me what you come out to get.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):